OpenStudy (anonymous):

simplify 4(1-6y-9y^2)=7(1-3y)+1

OpenStudy (anonymous):

4(1-6y-9y^2)=1+7(1-3y)

OpenStudy (anonymous):

4+4(-1)6y + 4(-1)9y^2 = 1+7+7(-1)3y

OpenStudy (anonymous):

can you do it now?